If fourteen cards are dealt from the top of a well shuffled deck of cards one after another without replacement, what is the chance that the top card is a heart?
Answer #1

here as there are 13 cards of hearts out of total 52 cards ; and each card has equal chance to be on top

therefore chance that the top card is a heart =13/52 =1/4